# Offline Model-Based Optimization: Comprehensive Review

### v1

@@ 3.1.1 Synthetic Function
(1) Computational efficiency: With closed-form expressions, these functions are quick to compute, allowing for extensive experimentation. (2) Scalability: Their ability to be defined for arbitrary input dimensions and numbers of objectives makes them highly adaptable for evaluating algorithms on large-scale and many-objective problems. (3) Analytical transparency: The known analytical forms enable exact computation of the Pareto front along with other key properties like gradients and Hessians, which helps in thoroughly understanding both the problem landscape and the behavior of optimization algorithms.
@@
Below, we introduce specific synthetic function tasks for both single-objective and multi-objective optimization, covering most tasks from commonly used benchmarks such as the BayesO benchmark (Kim, 2023; Surjanovic & Bingham, 2013), holo-bench (Stanton et al., 2024), pymoo (Blank & Deb, 2020), and Off-MOO-Bench (Xue et al., 2024). We describe each synthetic function class in Table 1. Due to the nature of synthetic function, the evaluation oracle is analytical, meaning that the mathematical form is known. Furthermore, the offline dataset size can be arbitrarily large (Any); however, researchers typically use defaults such as 50 0005000050\,00050 000 (Chen et al., 2024) for SOO or 60 0006000060\,00060 000 for MOO (Xue et al., 2024; Yuan et al., 2024a) (as indicated in parentheses). Similarly, many synthetic functions offer the flexibility to operate over any number of dimensions or objectives, and conventional practice often employs specific values (also indicated in parentheses). Finally, the design space in synthetic functions is generally continuous.
@@
  • Many local minima Functions with many local minima are designed to challenge an algorithm’s exploration capabilities by presenting a rugged landscape filled with numerous suboptimal traps. Common examples in this category include Ackley (Bäck, 1996), Bukin 6 (Jamil & Yang, 2013), Cosines (Kim & Choi, 2023), Drop-Wave (Jamil & Yang, 2013), Eggholder (Jamil & Yang, 2013), Gramacy & Lee, Griewank (Gramacy & Lee, 2010), Holder Table (Jamil & Yang, 2013), Kim1 (Kim & Choi, 2023), Levy (Jamil & Yang, 2013), Rastrigin (Jamil & Yang, 2013), and Shubert (Jamil & Yang, 2013).
  • Bowl-shaped Bowl-shaped functions with a convex landscape offer a comparatively smooth and unimodal environment. Their gradual curvature makes them ideal for assessing the convergence speed and stability of an algorithm under benign conditions. Representative functions in this class include Bohachevsky (Jamil & Yang, 2013), Kim2 (Kim & Choi, 2023), Kim3 (Kim & Choi, 2023), and Sphere (Picheny et al., 2013).
  • Plate-shaped Plate-shaped functions are characterized by expansive flat regions or plateaus where gradients are nearly zero over large areas of the search space. Although dedicated plate-shaped functions are less prevalent in standard benchmarks, some tasks like Zakharov (Jamil & Yang, 2013) exhibit plateau-like regions.
  • Valley-shaped Valley-shaped functions possess a narrow, elongated channel leading to the global optimum. This class—exemplified by the Three-Hump Camel (Jamil & Yang, 2013), Six-Hump Camel (Jamil & Yang, 2013), and Rosenbrock (Picheny et al., 2013)—tests an algorithm’s precision and its ability to efficiently traverse an ill-conditioned, elongated basin without overshooting the target.
  • Steep ridges/drops Functions characterized by steep ridges or abrupt drops exhibit sudden, dramatic changes in the objective landscape. Representative examples include De Jong 5 (Jamil & Yang, 2013), Easom (Jamil & Yang, 2013), and Michalewicz (Jamil & Yang, 2013).
  • Others There are some synthetic benchmark functions that do not neatly fit into the previously discussed classes but are nonetheless widely used in single-objective optimization. Examples include Beale (Jamil & Yang, 2013), Branin (Picheny et al., 2013), Colville (Jamil & Yang, 2013), Goldstein-Price (Picheny et al., 2013), Hartmann 3D (Jamil & Yang, 2013), and Hartmann 6D (Jamil & Yang, 2013). Each of them exhibits unique landscape features such as nonlinearity, moderate modality, or intricate curvature.
    Table 1: Overview of the synthetic function tasks.
    Task Name Size Space Dimension # Obj Oracle
    Many local minima Any (50000) Continuous 1 - Any (60) 1 Analytical
    Bowl-shaped Any (50000) Continuous 2 - Any (60) 1 Analytical
    Plate-shaped Any (50000) Continuous Any (60) 1 Analytical
    Valley-shaped Any (50000) Continuous 2 - Any (60) 1 Analytical
    Steep ridges/drops Any (50000) Continuous 2 1 Analytical
    Other SOO tasks Any (50000) Continuous 2 - 6 1 Analytical
    DTLZ Any (60000) Continuous Any (7 - 10) Any (3) Analytical
    Omnitest Any (60000) Continuous 2 2 Analytical
    VLMOP Any (60000) Continuous 1 - 6 2 - 3 Analytical
    ZDT Any (60000) Continuous 10 - 30 2 Analytical
@@
In addition to SOO tasks, many synthetic functions have been developed for the MOO setting, where algorithms often optimize multiple conflicting objectives simultaneously. We discuss four families of synthetic functions for MOO—DTLZ, Omnitest, VLMOP, and ZDT—each characterized by distinct Pareto front properties and varying levels of complexity.
  • DTLZ The DTLZ family (Deb et al., 2005) is designed to be scalable with respect to both the number of objectives and the decision variables, making them versatile for testing algorithms under various conditions. For example, DTLZ1 features a linear Pareto front, while DTLZ2 through DTLZ7 introduce different degrees of nonlinearity, multimodality, and deceptive features.
  • Omnitest Omnitest typically involves two objectives and presents a challenging landscape marked by convexity and multiple local optima (Deb & Tiwari, 2008).
  • VLMOP These functions (van Veldhuizen & Lamont, 1999) are characterized by intricate, concave or disconnected Pareto fronts, which challenge algorithms to capture a diverse set of trade-off solutions.
  • ZDT Functions in the ZDT family exhibit diverse Pareto front geometries, from convex to concave and even disconnected fronts (Zitzler et al., 2000). For example, ZDT1 and ZDT4 feature convex Pareto fronts, whereas ZDT2 and ZDT6 exhibit concave ones. Unlike other functions in the ZDT family, ZDT3 has a disconnected Pareto front.
@@ 3.1.2 Real-World System
Real-world system tasks encapsulate the inherent complexity of practical engineering design. While many of these tasks still utilize analytical functions for construction, unlike the purely synthetic functions, they are derived from models of real-world systems (Tanabe & Ishibuchi, 2020). Alternatively, some tasks employ simulations (Brockman et al., 2016) or surrogate models (Tanabe & Ishibuchi, 2020) as oracles. Overall, compared to the synthetic function category, these tasks incur higher evaluation costs and provide less transparency into the underlying mechanisms.
@@
The industrial suite benchmarks (Qian et al., 2025) further enrich this category, which are formulated as constrained optimization problems. For instance, the optimal operation task targets improvements in chemical processing quality, where the objective is to enhance the alkylating product subject to $14$ constraints designed to limit onboard fuel and launcher performance. The process flow sheeting problem and process synthesis problem focus on efficient process design under practical constraints. Both objective functions and constraints are expressed mathematically for these tasks. Besides, some classic combinatorial optimization problems in industrial applications—such as the multi-objective traveling salesman problem (Lust & Teghem, 2010), the multi-objective capacitated vehicle routing problem (Zajac & Huber, 2021a), and the multi-objective knapsack problem (Ishibuchi et al., 2015)—have analytical formulations that allow for rapid computation. However, these problems are not typically considered traditional design problems.
@@
Sometimes, even when the underlying mechanisms of a task are well understood, their complexity prevents us from expressing the oracle in a closed analytical form. In such cases, we rely on numerical simulations to obtain the oracle, despite the high computational cost often involved. A typical example is robot morphology design. Robot morphology design tasks focus on optimizing the structural parameters of simulated robots to maximize performance on specific tasks, as exemplified by the Ant and D’Kitty Morphology tasks (Trabucco et al., 2022). In the Ant Morphology task, the objective is to optimize $60$ continuous parameters—encoding limb size, orientation, and location—for a quadruped robot from OpenAI Gym (Brockman et al., 2016) so that it runs as fast as possible. In the D’Kitty Morphology task, the goal is to adjust $56$ continuous parameters defining the D’Kitty robot from ROBEL (Ahn et al., 2020) to enable a pre-trained, morphology-conditioned neural network controller, optimized using Soft Actor Critic (Haarnoja et al., 2018), to navigate the robot to a fixed location. Both tasks utilize the MuJoCo simulation (Todorov et al., 2012) to run simulations for $100$ time steps and average the results over $16$ independent trials, thereby providing reliable yet computationally efficient performance estimates.
@@
Table 2: Overview of the real-world system tasks.
Task Name Size Space Dimension # Obj Oracle
RE suite Any (60000) Mixed 3 - 7 2 - 9 Analytical / Surrogate
Industrial suite Any (2000 - 7000) Continuous 2 - 7 1 Analytical
Ant / D’Kitty 25009 Continuous 60 / 56 1 Simulation
Electronics design 8000 Categorical 10 1 Simulation
@@ 3.1.3 Scientific Design
Below, we present specific scientific design tasks across various domains—namely biology, chemistry, and materials science. As summarized in Table 3, these benchmarks offer diverse and challenging scenarios for offline MBO, reflecting the complexity of real-world scientific problems. Given the inherent nature of scientific design, the design space is typically discrete, and the evaluation oracles are implemented via surrogate, simulation, analytical, or lookup table.
@@
    (1) Green Fluorescent Protein (GFP): Optimizes a $237$-length protein sequence to increase fluorescence, using a dataset of $56,086$ variants (Sarkisyan et al., 2016a). (2) 5’ Untranslated Region (UTR) : Designs a $50$-nucleotide sequence to maximize gene expression, predicting ribosome load based on $280,000$ sequences (Sample et al., 2019a). (3) Adeno-Associated Virus (AAV): Optimizes a $28$-amino acid segment of the VP1 protein using $284,000$ variants, targeting improved viral viability (Ogden et al., 2019). (4) E4B Ubiquitination Factor (E4B): Enhances ubiquitination activity by optimizing $100,000$ mutations (Starita et al., 2013). (5) TEM-1 $\beta$-Lactamase (TEM): Aims to improve thermodynamic stability, leveraging $17,857$ sequences (Ren et al., 2022). (6) Aliphatic Amide Hydrolase (AMIE): Focuses on boosting enzyme activity, using $6,629$ variants (Wrenbeck et al., 2017). (7) Levoglucosan Kinase (LGK): Targets enzyme performance improvements, based on $7,891$ mutants (Klesmith et al., 2015). (8) Poly(A)-Binding Protein 1 (Pab1): Optimizes RNA binding efficiency with over $36,000$ sequences (Melamed et al., 2013). (9) SUMO E2 Conjugase (UBE2I): Focuses on protein function optimization, based on $2,000$ variants (Weile et al., 2017). (10) Regex: Modifies sequences to maximize bigram counts, simulating sequence editing operations (Stanton et al., 2022). (11) Affinity maturation (Affinity): Mutates antibody to maximize binding affinity with the antigen (Chen et al., 2025). (12) Red Fluorescent Protein (RFP): A multi-objective optimization problem balancing stability and solvent-accessible surface area (Stanton et al., 2022).
@@
  • DNA DNA design involves the optimization of sequences comprising four nucleotides (A, C, G, and T). (1) Transcription Factor Binding 8 (TFB8) : Optimizes transcription factor binding activity in an $8$-length sequence space (Barrera et al., 2016a). (2) Transcription Factor Binding 10 (TFB10): Extends the optimization to a $10$-length sequence space, aiming for higher binding affinity (Barrera et al., 2016a). Trabucco et al. (2022) further process the data to ensure training set given to offline MBO methods is restricted to the bottom $50\%$. Oracle evaluation is performed via a lookup table oracle, as the properties of all possible sequences have been pre-measured.
@@
Table 3: Overview of the scientific design tasks.
Task Name Size Space Dimension # Obj Oracle
GFP 56086 Categorical 237 1 Surrogate
UTR 280000 Categorical 50 1 Surrogate
AAV 284000 Categorical 28 1 Surrogate
E4B 100000 Categorical 102 1 Surrogate
TEM 17857 Categorical 286 1 Surrogate
AMIE 6629 Categorical 341 1 Surrogate
LGK 7891 Categorical 439 1 Surrogate
Pab1 36000 Categorical 75 1 Surrogate
UBE2I 2000 Categorical 159 1 Surrogate
Regex 42048 Categorical 73 3 Analytical
Affinity 4158 Categorical N/A 1 Simulation
RFP 4937 Categorical 489 2 Simulation
TFB8 32898 Categorical 8 1 Lookup Table
TFB10 50000 Categorical 10 1 Lookup Table
RNA-Binding 5000 Categorical 14 1 Simulation
Ehrlich Any (50000) Continuous 7 1 Analytical
ChEMBL 1093 Categorical 31 1 Surrogate
ZINC 48000 Categorical 257 2 Analytical
Molecule 49001 Continuous 32 2 Surrogate
Superconductor 21263 Continuous 86 1 Surrogate
@@
Unlike biological tasks focusing on large molecules such as proteins, and chemical tasks centering on small molecules, the material science category is dedicated to designing and optimizing materials based on complex compositional and structural performance metrics. A typical offline MBO benchmark in this domain is Superconductor. Derived from a dataset of $21{,}263$ superconductors with annotated critical temperatures (Hamidieh, 2018), it has also been incorporated into design-bench (Trabucco et al., 2022). The goal is to maximize the critical temperature by optimizing an $86$-dimensional continuous representation of material composition, with a random forest surrogate serving as the evaluation oracle.
@@
  • Battery materials design, where research efforts focus on discovering better active materials for lithium-ion batteries (Valladares & Others, 2021).
@@ 3.1.4 Machine Learning Model
Table 4 summarizes key characteristics of these tasks, covering common benchmarks used in offline MBO research. We categorize these tasks into three major subcategories—model architecture, model parameter, and model hyperparameter—which are described in detail below.
@@
Table 4: Overview of the machine learning model tasks.
Task Name Size Space Dimension # Obj Oracle
NAS 1771 Categorical 64 1 Real Experiment
NAS-Bench-201-Test 9375 Categorical 6 3 Lookup Table
C-10/MOP 9375 - 60000 Categorical 5 - 32 2 - 8 Lookup Table
IN-1K/MOP 60000 Categorical 21 - 34 2 - 4 Lookup Table
Hopper 3200 Continuous 5126 1 Simulation
MO-Swimmer 8571 Continuous 9734 2 Simulation
MO-Hopper 4500 Continuous 10184 2 Simulation
HPOBench 10000 Mixed 2-27 1 Lookup Table / Surrogate
@@ 3.2.1 Usefulness
Since the true Pareto front is generally unknown in real-world tasks, most existing studies primarily rely on HV to evaluate MOO performance (Tanabe & Ishibuchi, 2020; Xue et al., 2024; Yuan et al., 2024a).
@@ 3.2.2 Diversity
Diversity quantifies the spread or variability within the set of generated candidate designs (Jain et al., 2022; Kim et al., 2023; Kirjner et al., 2024). In offline MBO, it is crucial not only to identify high-performing designs but also to explore diverse regions of the design space, thereby capturing multiple modes of the black-box function. A common metric for measuring diversity is the average pairwise distance between all distinct candidates in the set $\mathcal{B}$, defined as:
@@
[…]
@@ 3.2.3 Novelty
Novelty measures the degree to which the newly generated candidate set $\mathcal{B}$ differs from the offline dataset $\mathcal{D}$. In offline MBO, it is important that the proposed candidates not only perform well but also explore regions of the design space that are distinct from known designs. Novelty is typically quantified by computing the mean of the minimum distances between each candidate in $\mathcal{B}$ and the closest design in $\mathcal{D}$ (Jain et al., 2022; Kim et al., 2023). Mathematically, the novelty score is defined as:

### v2

@@ 3.1.1 Synthetic Function
(1) Computational efficiency: With closed-form expressions, these functions are quick to compute, allowing for extensive experimentation. (2) Scalability: Their ability to be defined for arbitrary input dimensions and numbers of objectives makes them highly adaptable for evaluating algorithms on large-scale problems. (3) Analytical transparency: The known analytical forms enable exact computation of the Pareto front along with other key properties like gradients and Hessians, which helps in thoroughly understanding both the problem landscape and the behavior of optimization algorithms.
@@
Below, we introduce specific synthetic function tasks for both single-objective and multi-objective optimization, covering most tasks from commonly used benchmarks such as the BayesO benchmark (Kim, 2023; Surjanovic & Bingham, 2013), holo-bench (Stanton et al., 2024), pymoo (Blank & Deb, 2020), and Off-MOO-Bench (Xue et al., 2024). Due to the nature of synthetic function, the evaluation oracle is analytical, meaning that the mathematical form is known. Furthermore, the offline dataset size can be arbitrarily large (Any); however, researchers typically use defaults such as 50 00050\,000 (Chen et al., 2024) for SOO or 60 00060\,000 for MOO (Xue et al., 2024; Yuan et al., 2025). Similarly, many synthetic functions offer the flexibility to operate over any number of dimensions or objectives, and conventional practice often employs specific values (also indicated in parentheses). Finally, the design space in synthetic functions is generally continuous.
@@
In addition to SOO tasks, many synthetic functions have been developed for the MOO setting, where algorithms often optimize multiple conflicting objectives simultaneously. We discuss four families of synthetic functions for MOO—DTLZ (Deb et al., 2005), Omnitest (Deb & Tiwari, 2008), VLMOP (van Veldhuizen & Lamont, 1999), and ZDT (Zitzler et al., 2000)—each characterized by distinct Pareto front properties and varying levels of complexity.
@@ 3.1.2 Real-World System
Real-world system tasks encapsulate the inherent complexity of practical engineering design. The underlying mechanisms of these designs are generally well-understood, in contrast to the scientific design tasks discussed in the next subsection. While many of these tasks still utilize analytical functions for construction, unlike the purely synthetic functions, they are derived from models of real-world systems (Tanabe & Ishibuchi, 2020). Alternatively, some tasks employ simulations (Brockman et al., 2016) or surrogate models (Tanabe & Ishibuchi, 2020) as oracles. Overall, compared to the synthetic function category, these tasks incur higher evaluation costs and provide less transparency into the underlying mechanisms.
@@
The industrial suite benchmarks (Qian et al., 2025) further enrich this category, which are formulated as constrained optimization problems. For instance, the optimal operation task targets improvements in chemical processing quality, where the objective is to enhance the alkylating product subject to $14$ constraints designed to limit onboard fuel and launcher performance. The process flow sheeting problem and process synthesis problem focus on efficient process design under practical constraints. Both objective functions and constraints are expressed mathematically for these tasks. Besides, some classic combinatorial optimization problems in industrial applications—such as the multi-objective traveling salesman problem (Lust & Teghem, 2010), the multi-objective capacitated vehicle routing problem (Zajac & Huber, 2021), and the multi-objective knapsack problem (Ishibuchi et al., 2015)—have analytical formulations that allow for rapid computation. However, these problems are not typically considered traditional design problems.
@@
Sometimes, even when the underlying mechanisms of a task are well understood, their complexity prevents us from expressing the oracle in a closed analytical form. In such cases, numerical simulations are used to obtain the oracle, despite the high computational cost often involved. A typical example is robot morphology design. Robot morphology design tasks focus on optimizing the structural parameters of simulated robots to maximize performance on specific tasks, as exemplified by the Ant and D’Kitty Morphology tasks (Trabucco et al., 2022). In the Ant Morphology task, the objective is to optimize $60$ continuous parameters—encoding limb size, orientation, and location—for a quadruped robot from OpenAI Gym (Brockman et al., 2016) so that it runs as fast as possible. In the D’Kitty Morphology task, the goal is to adjust $56$ continuous parameters defining the D’Kitty robot from ROBEL (Ahn et al., 2020) to enable a pre-trained, morphology-conditioned neural network controller, optimized using Soft Actor Critic (Haarnoja et al., 2018), to navigate the robot to a fixed location. Both tasks utilize the MuJoCo simulation (Todorov et al., 2012) to run simulations for $100$ time steps and average the results over $16$ independent trials, thereby providing reliable yet computationally efficient performance estimates.
@@ 3.1.3 Scientific Design
Below, we present specific scientific design tasks across various domains—namely biology, chemistry, and materials science. These benchmarks offer diverse and challenging scenarios for offline MBO, reflecting the complexity of scientific problems. Given the inherent nature of scientific design, the design space is typically discrete, and evaluation oracles are implemented via surrogate, simulation, analytical, or lookup table.
@@
    (1) Green Fluorescent Protein (GFP): Optimizes a $237$-length protein sequence to increase fluorescence, using a dataset of $56,086$ variants (Sarkisyan et al., 2016). (2) 5’ Untranslated Region (UTR) : Designs a $50$-nucleotide sequence to maximize gene expression, predicting ribosome load based on $280,000$ sequences (Sample et al., 2019). (3) Adeno-Associated Virus (AAV): Optimizes a $28$-amino acid segment of the VP1 protein using $284,000$ variants, targeting improved viral viability (Ogden et al., 2019). (4) E4B Ubiquitination Factor (E4B): Enhances ubiquitination activity by optimizing $100,000$ mutations (Starita et al., 2013). (5) TEM-1 $\beta$-Lactamase (TEM): Aims to improve thermodynamic stability, leveraging $17,857$ sequences (Ren et al., 2022). (6) Aliphatic Amide Hydrolase (AMIE): Focuses on boosting enzyme activity, using $6,629$ variants (Wrenbeck et al., 2017). (7) Levoglucosan Kinase (LGK): Targets enzyme performance improvements, based on $7,891$ mutants (Klesmith et al., 2015). (8) Poly(A)-Binding Protein 1 (Pab1): Optimizes RNA binding efficiency with over $36,000$ sequences (Melamed et al., 2013). (9) SUMO E2 Conjugase (UBE2I): Focuses on protein function optimization, based on $2,000$ variants (Weile et al., 2017). (10) Regex: Modifies sequences to maximize bigram counts, simulating sequence editing operations (Stanton et al., 2022). (11) Affinity maturation (Affinity): Mutates antibody to maximize binding affinity with the antigen (Chen et al., 2025). (12) Red Fluorescent Protein (RFP): A multi-objective optimization problem balancing stability and solvent-accessible surface area (Stanton et al., 2022).
@@
  • DNA DNA design involves the optimization of sequences comprising four nucleotides (A, C, G, and T). (1) Transcription Factor Binding 8 (TFB8) : Optimizes transcription factor binding activity in an $8$-length sequence space (Barrera et al., 2016). (2) Transcription Factor Binding 10 (TFB10): Extends the optimization to a $10$-length sequence space, aiming for higher binding affinity (Barrera et al., 2016). Trabucco et al. (2022) further process the data to ensure training set given to offline MBO methods is restricted to the bottom $50\%$. Oracle evaluation is performed via a lookup table oracle, as the properties of all possible sequences have been pre-measured.
@@
Unlike biological tasks focusing on large molecules such as proteins, and chemical tasks centering on small molecules, the material science category is dedicated to designing and optimizing materials based on complex compositional and structural performance metrics. A typical offline MBO benchmark in this domain is Superconductor. Derived from a dataset of 21,26321{,}263 superconductors with annotated critical temperatures (Hamidieh, 2018), it has also been incorporated into design-bench (Trabucco et al., 2022). The goal is to maximize the critical temperature by optimizing an $86$-dimensional continuous representation of material composition, with a random forest surrogate serving as the evaluation oracle.
@@
  • Battery materials design, where research efforts focus on discovering better active materials for lithium-ion batteries (Valladares et al., 2021).
@@ 3.1.4 Machine Learning Model
We categorize these tasks into three major subcategories—model architecture, model parameter, and model hyperparameter—which are described in detail below.
@@ 3.2.1 Usefulness
Since the true Pareto front is generally unknown in real-world tasks, most existing studies primarily rely on HV to evaluate MOO performance (Tanabe & Ishibuchi, 2020; Xue et al., 2024; Yuan et al., 2025).
@@ 3.2.2 Diversity
Diversity quantifies the spread or variability within the set of generated candidate designs, and maintaining higher diversity helps reduce the risk that all candidates fail due to similar underlying limitations (Jain et al., 2022; Kim et al., 2023; Kirjner et al., 2024). In offline MBO, it is crucial not only to identify high-performing designs but also to explore diverse regions of the design space, thereby capturing multiple modes of the black-box function. A common metric for measuring diversity is the average pairwise distance between all distinct candidates in the set $\mathcal{B}$, defined as:
@@
$\mathrm{L1C}\bigl(\mathcal{B}\bigr)=\frac{1}{d}\sum_{k=1}^{d}\max_{i\neq j}\Bigl\lvert x_{ik}-x_{jk}\Bigr\rvert$,
@@ 3.2.3 Novelty
Novelty measures the degree to which the newly generated candidate set $\mathcal{B}$ differs from the offline dataset $\mathcal{D}$, thus promoting exploration beyond known designs.. In offline MBO, it is important that the proposed candidates not only perform well but also explore regions of the design space that are distinct from known designs. Novelty is typically quantified by computing the mean of the minimum distances between each candidate in $\mathcal{B}$ and the closest design in $\mathcal{D}$ (Jain et al., 2022; Kim et al., 2023). Mathematically, the novelty score is defined as: